# Equivariant Neural Networks for General Linear Symmetries on Lie Algebras

### v1

@@ 4.1 A General $\mathrm{Ad}$-invariant Bilinear Form for Reductive Lie Algebras
where $B$ denotes the Killing form on $\mathfrak{g},\mathfrak{g}]$.
Proposition 4.1.
The bilinear form $\widetilde{B}$ is symmetric, $\mathrm{Ad}$-invariant, and nondegenerate. Moreover, $\mathfrak{z}(\mathfrak{g})$ and $\mathfrak{g},\mathfrak{g}]$ are $\widetilde{B}-$orthogonal, with $\widetilde{B}|_{[\mathfrak{g},\mathfrak{g}]}=B$ and $\widetilde{B}|_{\mathfrak{z}(\mathfrak{g})}=\langle\cdot,\cdot\rangle_{\mathfrak{z}}$.
@@
$B$ vanishes on $\mathfrak{z}(\mathfrak{g})$ and is $\mathrm{Ad}$-invariant; by construction $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ is $\mathrm{Ad}$-invariant. Symmetry is immediate. Nondegeneracy follows since $B$ is nondegenerate on the semisimple ideal and $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ is nondegenerate on the center; orthogonality holds because $B(\mathfrak{z}(\mathfrak{g}),[\mathfrak{g},\mathfrak{g}])=0$. ∎
For our primary case $\mathfrak{g}=\mathfrak{gl}(n,\mathbb{R})=\mathbb{R}I\oplus\mathfrak{sl}(n,\mathbb{R})$ we choose the canonical trace-based form
$\widetilde{B}(X,Y)\;=\;2n\cdot\mathrm{tr}(XY)\;-\;\mathrm{tr}(X)\,\mathrm{tr}(Y),$ (2)
This form $\widetilde{B}$ is the fundamental tool used throughout the ReLN architecture.
@@
Our concrete form in Eq. 2 satisfies the conditions of Proposition 4.1. Decomposing a matrix $X=X_{0}+\frac{1}{n}\mathrm{tr}(X)I$ (where $X_{0}\in\mathfrak{sl}(n)$) reveals that our form separates orthogonally:
@@
This decomposition directly shows how $\widetilde{B}$ serves as a generalization of prior work. The first term is the Killing form on the semisimple part, which is the tool used in Lie Neurons (Lin et al., 2024a). The second term is a standard inner product on the center, which, under the isomorphism $\mathfrak{so}(3)\simeq\mathbb{R}^{3}$, recovers the dot product used in Vector Neurons (Deng et al., 2021). Our single form thus unifies these approaches, extending to the full reductive algebra $\mathfrak{gl}(n)$ (details can be found in Appendix C).
@@ 4.2 The ReLN Layer Toolbox
A linear map applied to the channel dimension $f(x;W)=xW$ with $W\in\mathbb{R}^{C\times C^{\prime}}$ is strictly equivariant: the group acts on the left (geometric dimension) while $W$ acts on the right (channel dimension), and thus these operations commute (formal proof in Appendix D).
@@
We introduce two complementary nonlinear primitives; full definitions, parameterizations, and stability prescriptions are deferred to Appendix D.
  • ReLN-ReLU: To overcome the non-equivariance of elementwise activations, we use our form $\widetilde{B}$ to define a directional nonlinearity. Each vector feature $x_{c}$ is rectified along a learnable direction $d_{c}$ via update rule:
  • ReLN-ReLU Using $\widetilde{B}$ to build an invariant gate, each channel feature is updated by
    $x_{c}^{\prime}\;=\;x_{c}\;+\;\max\big(0,\;\widetilde{B}(x_{c}^{\wedge},d_{c}^{\wedge})\big)\,d_{c},$ (4)
@@
  • Max-Killing Pooling: To aggregate a set of features $\{X_{n}\}_{n=1}^{N}$, where each $X_{n}$ is a multi-channel feature tensor, this layer selects the representative feature with the maximal projection onto a learnable direction. For each channel $c$, the index is found via $n^{*}(c)=\arg\max_{n}\widetilde{B}(X_{n,c},D_{n,c})$, and the pooled feature is $X^{\max}_{c}=X_{n^{*}(c),c}$.
@@ 5 Experiments
We evaluate ReLNs on a suite of tasks designed to highlight two complementary strengths: algebraic generality on benchmarks and practical efficacy on a challenging, uncertainty-aware robotics task. We compare against standard non-equivariant baselines (MLP, ResNet), the original Lie Neurons, and a Vector Neurons–style baseline adapted for covariance inputs by eigendecomposition. For the Top-Tagging benchmark we also report results versus established physics models used in prior work.
@@ 5.1.1 Platonic Solid Classification
We first validate our model on the Platonic solid classification benchmark from (Lin et al., 2024a), testing the adjoint-equivariance where camera rotations induce a conjugation action on inter-face homographies in $\mathrm{SL}(3)$. Full experimental details are provided in Appendix E.2.
The results, summarized in Table 2, confirm that non-equivariant baselines fail to generalize to rotated camera views. This fundamental weakness is not resolved by data augmentation or increased model capacity, as our wider MLP variant with approximately double the parameters shows negligible improvement on the out-of-distribution test set. In contrast, the ReLN model achieves near-perfect accuracy with robustness, matching the performance of the Lie Neurons while demonstrating improved results. Importantly, this result validates that our general $\mathfrak{gl}(n)$ framework operates effectively on common semisimple subalgebras, as its built-in adjoint-equivariance on the parent group yields robust behavior when restricted to subgroups like $\mathrm{SO}(3)$ and $\mathrm{SL}(3)$.
@@
To further probe our framework’s algebraic generality, our second benchmark involves regressing a highly nonlinear invariant function on the real symplectic Lie algebra $\mathfrak{sp}(4,\mathbb{R})$. The symplectic algebra $\mathfrak{sp}(2n)$ is the mathematical foundation of Hamiltonian mechanics, which describes any physical system where energy is conserved. Our target is a scalar invariant defined for pairs $X,Y\in\mathfrak{sp}(4,\mathbb{R})$:
$g(X,Y)=\sin\!\big(\operatorname{Tr}(XY)\big)+\cos\!\big(\operatorname{Tr}(YY)\big)-\tfrac{1}{2}\operatorname{Tr}(YY)^{3}+\det(XY)+\exp\!\big(\operatorname{Tr}(XX)\big).$ (6)
We generate a dataset of 10k training and 10k test pairs by sampling from $\mathfrak{sp}(4,\mathbb{R})$. We compare ReLN against MLP baselines (trained with and without $\mathrm{Sp}(4)$ data augmentation) and the original Lie Neurons. At test time, we report the standard MSE, MSE averaged over 500 random adjoint actions, and the invariance error.
As shown in Table 3, non-equivariant MLPs are orders of magnitude less accurate and exhibit high invariance error, failing to learn the group structure even with data augmentation. Our ReLN model not only achieves the lowest MSE and near-zero invariance error, but also shows a modest but consistent improvement over Lie Neurons. This suggests that our non-degenerate bilinear form provides not only theoretical generality but also superior numerical conditioning in practice.
Table 3: Regression performance and invariance error on $\mathfrak{sp}(4)$. “Training Aug.” indicates whether $\mathrm{Sp}(4)$ conjugation was applied during training.
Model Training Aug. # Params Test Aug. Invariance Error
@@
5.2 Particle Physics with Lorentz Group $\mathrm{SO}(1,3)$ Equivariance
We test our framework on the Top-Tagging benchmark (Kasieczka et al., 2019), a task to distinguish particle jets originating from top quarks against a large background from Quantum Chromodynamics. Because these relativistic collisions are subject to the symmetries of spacetime, the task requires equivariance under the Lorentz group $\mathrm{SO}(1,3)$. We solve this left-action equivariant learning problem by introducing a map that embeds the four-momentum $p\in\mathbb{R}^{1,3}$ of each constituent particle, which combines its energy and 3D momentum, into the Lie algebra $\mathfrak{gl}(5)$ as $\varphi(p)=\begin{pmatrix}0_{4\times 4}&p\\ p^{\top}\eta&0\end{pmatrix}$, where $\eta=\mathrm{diag}(-1,1,1,1)$ is the Minkowski metric. As proven in Appendix F.1, this embedding unifies left- and adjoint-action equivariance within a single Lie-algebraic framework. We adapt the LorentzNet architecture by replacing its invariant feature computation with our proposed bilinear form. To ensure a fair comparison, we created a parameter-matched version of LorentzNet. As shown in Table 4, our model achieves competitive overall performance while demonstrating an advantage on the background rejection metric. This result shows that our general Lie-algebraic approach can effectively unify adjoint- and left-action equivariance in a parameter-efficient manner.
Table 4: Comparison of performance on the Top-Tagging dataset. Rej@30% denotes the background rejection at 30% signal efficiency (higher is better). Benchmark scores are as reported in the original publications.
Architecture #Params Accuracy AUC Rej@30% Reference
PELICAN 45k 0.943 0.987 $2289\pm 204$ Bogatskiy et al. (2022)
LorentzNet (original) 224k 0.942 0.987 $2195\pm 173$ Gong et al. (2022)
LorentzNet (param-matched) 84k 0.942 0.987 $1821\pm 94$ Our reproduction
LGN 4.5k 0.929 0.964 $435\pm 95$ Bogatskiy et al. (2020)
BIP 4k 0.931 0.981 $853\pm 68$ Munoz et al. (2022)
partT 2.14M 0.940 0.986 $1602\pm 81$ Qu et al. (2022)
ParticleNet 498k 0.938 0.985 $1298\pm 46$ Qu & Gouskos (2020)
EFN 82k 0.927 0.979 $633\pm 31$ Komiske et al. (2019)
TopoDNN 59k 0.916 0.972 $295\pm 5$ Pearkes et al. (2017)
LorentzMACE 228k 0.942 0.987 $1935\pm 85$ Batatia et al. (2023)
ReLN (Ours) 84k 0.942 0.987 $1979\pm 87$
@@ 5.3 Drone State Estimation with Geometric Uncertainty
We test our framework on a challenging drone state estimation task using a large-scale dataset of aggressive, highly dynamic flights. The objective is to recover a 3D trajectory from a sequence of noisy velocity measurements and their corresponding time-varying covariances, where each covariance matrix quantifies the uncertainty of its associated velocity measurement. This setup tests a model’s ability to jointly process vector ($\mathbf{v}$) and matrix ($C$) data in a geometrically consistent and uncertainty-aware manner.
@@
We created a large-scale synthetic dataset of 200 aggressive drone trajectories, over 13 hours of challenging, high-speed flight (details in Appendix G). The network is trained to regress the 3D position from a sequence of velocity and covariance measurements within a time window. Our model is compared against non-equivariant ResNets and an $\mathrm{SO}(3)$-equivariant baseline using Vector Neurons (VN). Since VNs cannot directly process matrix inputs, we implement a method that handles covariance matrices via an eigendecomposition-based strategy (details in Appendix G.3).
In contrast, our ReLN architecture treats both velocity and covariance as unified geometric objects within a single algebraic space, $\mathfrak{gl}(3)$. Specifically, the velocity vector $\mathbf{v}\in\mathbb{R}^{3}$ is lifted to its matrix representation in $\mathfrak{so}(3)\subset\mathfrak{gl}(3)$. For the time-varying covariance $C$, we explore two representations that also reside in $\mathfrak{gl}(3)$: (1) the matrix $C$ directly, and (2) its matrix logarithm $\log C$, which respects the geometry of the $\mathrm{SPD}(3)$ manifold. This unified representation ensures a measurement and its uncertainty transform consistently under the adjoint action of $\mathrm{SO}(3)$. The final velocity estimate is then equivariantly extracted by projecting the network’s matrix output onto its skew-symmetric component. We test three ReLN variants: (1) velocity only, (2) velocity + covariance, and (3) velocity + log-covariance. Full implementation details are provided in Appendix G.
@@ 5.3.1 Results and Analysis
Table 5: Performance on the drone trajectory dataset. Best result in each column is shown in bold.
Model ID $\mathrm{SO}(3)$
ATE ATE% RPE ATE ATE% RPE
Non-Equivariant Baselines
   ResNet (Velocity only) $208.07$ $95.06$ $107.60$ $217.02$ $100.39$ $111.29$
   ResNet (Velocity + Covariance) $205.11$ $94.94$ $106.07$ $213.26$ $98.90$ $109.37$
Equivariant Baselines
   VN (Velocity only) $17.36$ $7.52$ $13.51$ $17.36$ $7.52$ $13.51$
   VN (Velocity + Covariance) $191.78$ $88.66$ $98.39$ $190.22$ $88.47$ $98.26$
@@
   ReLN (Velocity only) $16.85$ $7.31$ $12.7$ $16.85$ $7.31$ $12.7$
   ReLN (Velocity + Covariance) $16.49$ $7.21$ $13.02$ $16.49$ $7.21$ $13.02$
   ReLN (Velocity + $\log$-Covariance) 13.92 5.99 11.04 13.92 5.99 11.04
Abbreviations and units: ID = in-distribution test (no rotation), $\mathrm{SO}(3)$ = test-time random $\mathrm{SO}(3)$ rotations. ATE = absolute trajectory error (meters), ATE% = ATE relative to trajectory length (%), RPE = relative pose error (meters).

### v2

@@ 4.1 An $\mathrm{Ad}$-invariant bilinear form for reductive Lie algebras
where $B$ denotes the Killing form on the semisimple ideal $\mathfrak{g},\mathfrak{g}]$.
Proposition 4.2 (Non-degeneracy and $\mathrm{Ad}$-invariance).
The form $\widetilde{B}$ is symmetric, $\mathrm{Ad}$-invariant, and non-degenerate. Moreover, $\mathfrak{z}(\mathfrak{g})$ and $\mathfrak{g},\mathfrak{g}]$ are $\widetilde{B}-$orthogonal, with $\widetilde{B}|_{[\mathfrak{g},\mathfrak{g}]}=B$ and $\widetilde{B}|_{\mathfrak{z}(\mathfrak{g})}=\langle\cdot,\cdot\rangle_{\mathfrak{z}}$.
@@
$B$ vanishes on $\mathfrak{z}(\mathfrak{g})$ and is $\mathrm{Ad}$-invariant; by construction $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ is $\mathrm{Ad}$-invariant. Symmetry is immediate. Nondegeneracy follows since $B$ is nondegenerate on the semisimple ideal and $\langle\cdot,\cdot\rangle_{\mathfrak{z}}$ is nondegenerate on the center; orthogonality holds because $B(\mathfrak{z}(\mathfrak{g}),[\mathfrak{g},\mathfrak{g}])=0$. Detailed proofs are provided in Appendix C. ∎
For our primary case $\mathfrak{g}=\mathfrak{gl}(n)=\mathbb{R}I\oplus\mathfrak{sl}(n)$, we use the closed form
$\widetilde{B}(X,Y)\;=\;2n\,\operatorname{Tr}(XY)\;-\;\operatorname{Tr}(X)\operatorname{Tr}(Y).$ (2)
This $\widetilde{B}$ is our basic $\mathrm{Ad}$-invariant contraction throughout ReLN architecture.
@@
Our concrete form in Eq. 2 satisfies the conditions of Definition 4.1. Decomposing a matrix $X=X_{0}+\frac{1}{n}\mathrm{tr}(X)I$ (where $X_{0}\in\mathfrak{sl}(n)$) reveals that our form separates orthogonally:
@@
This decomposition directly shows how $\widetilde{B}$ serves as a generalization of prior work. The first term is the Killing form on the semisimple part, which is the tool used in Lie Neurons (Lin et al., 2024a). The second term is a standard inner product on the center, which, under the isomorphism $\mathfrak{so}(3)\simeq\mathbb{R}^{3}$, recovers the dot product used in Vector Neurons (Deng et al., 2021). For $\mathfrak{so}(3)$, the hat map intertwines the vector action and the adjoint action, so $\mathrm{Ad}$-equivariance on $\mathfrak{so}(3)$ corresponds to standard $\mathrm{SO}(3)$-equivariance on $\mathbb{R}^{3}$; see Lemma B.7. Our single form thus unifies these approaches, extending to the full reductive algebra $\mathfrak{gl}(n)$ (details can be found in Appendix B).
@@ 4.2 The ReLN Layer Toolbox
A linear map applied to the channel dimension $f(x;W)=xW$ with $W\in\mathbb{R}^{C\times C^{\prime}}$ is equivariant: the group acts on the left (geometric dimension) while $W$ acts on the right (channel dimension), and thus these operations commute (formal proof in Appendix E).
@@
We introduce two complementary nonlinear primitives; full definitions, parameterizations, and stability prescriptions are deferred to Appendix E.
  • ReLN-ReLU: To avoid basis-dependent elementwise activations, we use $\widetilde{B}$ to build an invariant scalar gate and apply it along an equivariant direction. For input features $x\in\mathbb{R}^{K\times C}$, we explicitly define the equivariant reference direction $d=xW$, where $W\in\mathbb{R}^{C\times C}$ is a learnable weight matrix. The feature is then updated via the following gated mechanism:
    $x_{c}^{\prime}=\begin{cases}x_{c},&\text{if }\widetilde{B}(x_{c}^{\wedge},d_{c}^{\wedge})\leq 0\\ x_{c}+\widetilde{B}(x_{c}^{\wedge},d_{c}^{\wedge})d_{c},&\text{otherwise}\end{cases}$ (4)
@@
[…]
@@ 5 Experiments
We evaluate ReLNs along three axes: (i) exact adjoint equivariance under broad linear symmetries (including reductive and non-compact groups), (ii) learning from coupled heterogeneous geometry (e.g., velocities with covariances), and (iii) accuracy–efficiency trade-offs against widely used equivariant and self-supervised baselines.
We consider (a) algebraic benchmarks for adjoint-equivariant learning on $\mathfrak{sl}(3)$ and $\mathfrak{sp}(4)$, (b) uncertainty-aware drone state estimation with joint velocity–covariance processing under $\mathrm{SO}(3)$ frame changes, (c) the EMLP double-pendulum benchmark assessing accuracy and per-step efficiency, and (d) learning from 3D Gaussian-splat primitives in ShapeSplat Ma et al. (2025) with comparisons to Gaussian-MAE. We compare against non-equivariant baselines (MLP, ResNet), Lie-algebraic equivariant models (Lie Neurons), and strong equivariant architectures including Vector Neurons (VN), Tensor Field Networks (TFN), and an $\mathrm{SE}(3)$-Transformer. Additional Lorentz-equivariant top-tagging results are deferred to Appendix H.
@@ 5.1.1 Platonic Solid Classification
We first validate our model on the Platonic solid classification benchmark from (Lin et al., 2024a), testing the adjoint-equivariance where camera rotations induce a conjugation action on the homographies between the projected faces of the solid, represented as $\mathrm{SL}(3)$. Full experimental details are provided in Appendix G.2.
The results, summarized in Table 1, confirm that non-equivariant baselines fail to generalize to rotated camera views. This generalization gap persists even with augmentation or increased capacity, as our wider MLP variant with approximately double the parameters shows negligible improvement on the out-of-distribution test set. In contrast, the ReLN model achieves near-perfect accuracy with robustness, matching the performance of the Lie Neurons while demonstrating improved results. This result validates that our general $\mathfrak{gl}(n)$ framework operates effectively on common semisimple subalgebras, as its built-in adjoint-equivariance on the parent group yields robust behavior when restricted to subgroups like $\mathrm{SO}(3)$ and $\mathrm{SL}(3)$.
@@
To probe algebraic generality beyond $\mathfrak{sl}(n)$, we regress a highly nonlinear $\mathrm{Sp}(4)$-invariant scalar on the real symplectic Lie algebra $\mathfrak{sp}(4,\mathbb{R})$ (the Lie algebra underlying Hamiltonian/symplectic symmetry). For $X,Y\in\mathfrak{sp}(4,\mathbb{R})$, the target is
$g(X,Y)=$ $\sin\!\big(\operatorname{Tr}(XY)\big)+\cos\!\big(\operatorname{Tr}(YY)\big)$
$-\tfrac{1}{2}\operatorname{Tr}(YY)^{3}+\det(XY)+\exp\!\big(\operatorname{Tr}(XX)\big).$ (6)
We sample 10k training and 10k test pairs and compare ReLN to MLP baselines (with/without $\mathrm{Sp}(4)$ conjugation augmentation) and Lie Neurons. We report test MSE, MSE averaged over 500 random adjoint actions, and the invariance error.
As shown in Table 2, non-equivariant MLPs are orders of magnitude less accurate and exhibit substantially larger invariance error, indicating that they fail to reliably capture the underlying group structure even with $\mathrm{Sp}(4)$ conjugation augmentation. In contrast, Lie Neurons and ReLN achieve low test MSE and near-zero invariance error, demonstrating stable invariant regression on $\mathfrak{sp}(4)$. At this parameter scale, the performance of the two equivariant models is comparable. These results support that our model matches the accuracy of specialized Lie-algebraic baselines while retaining a unified construction across diverse Lie algebras.
Table 2: Regression performance and invariance error on $\mathfrak{sp}(4)$. “Tr. Aug.” indicates training augmentation.
Model Tr. Aug. # Params Test Aug. Inv. Error
@@ 5.2 Drone State Estimation with Geometric Uncertainty
We evaluate ReLN on a drone state-estimation task that reconstructs 3D trajectories from noisy velocity measurements $\mathbf{v}\in\mathbb{R}^{3}$ and time-varying uncertainty covariances $C\in\mathrm{SPD}(3)$ during highly dynamic flights. This setup requires a model to jointly process vector and matrix data in a geometrically consistent and uncertainty-aware manner.
@@
We use a large-scale synthetic dataset of 200 aggressive trajectories (over 13 hours of flight); details are in Appendix I. We evaluate ReLN against: non-equivariant 1D ResNets, VN (Deng et al., 2021), and two spherical-harmonics-based equivariant architectures that use steerable bases and tensor-product coupling to construct equivariant messages: TFN (Thomas et al., 2018) and SE(3)-Transformers (Fuchs et al., 2020).
To isolate the effect of uncertainty fusion, we test three variants for each equivariant model: velocity-only, velocity + covariance, and velocity + log-covariance. For baselines such as VN that cannot natively process matrix inputs, we implement an eigendecomposition-based strategy to ingest covariance data (details in Appendix I.3). To ingest covariance data in TFN, SE(3)-Transformer, we apply irreducible representation (irrep) decomposition.
Irrep decomposition vs. unified adjoint processing.
Steerable baselines (TFN, SE(3)-Transformer) represent $C$ through irreducible components (trace and traceless-symmetric parts, corresponding to $\ell=0$ and $\ell=2$), and learn interactions with $\mathbf{v}$ via tensor-product coupling between fibers. This decomposition ignores the holistic geometry of the measurement-uncertainty pair, relying on tensor-product coupling to reconnect these components. See Appendix I.3 for the detailed architecture design and decomposition. In contrast, ReLN embeds both $\mathbf{v}$ and $C$ into a common matrix-valued representation with a single conjugation rule: $\mathbf{v}$ is lifted to $\mathfrak{so}(3)\subset\mathfrak{gl}(3)$ and $C$ (or $\log C$) is provided as a matrix feature. This yields a uniform $\mathrm{Ad}$-action for the pair and allows gating/normalization using the same $\mathrm{Ad}$-invariant non-degenerate form $\widetilde{B}$. The final velocity estimate is obtained by projecting the network output to its skew-symmetric component; implementation details are deferred to Appendix I.
@@ 5.2.1 Results and Analysis
Table 3: Drone state estimation performance (ATE and RTE in meters). $v,C,\log C$ denote velocity, covariance, and log-covariance inputs. Refer to Appendix I.6 for the exhaustive performance report including all variants and metrics.
Model Input ID SO(3)
ATE $\downarrow$ RTE $\downarrow$ ATE $\downarrow$ RTE $\downarrow$
Non-Equiv.
ResNet $(v,C)$ $205.11$ $106.07$ $213.26$ $109.37$
Equiv. Baselines
VN (Deng et al., 2021) $v$ $17.36$ $13.51$ $17.36$ $13.51$
VN (Deng et al., 2021) $(v,C)$ $191.78$ $98.39$ $190.22$ $98.26$
TFN (Thomas et al., 2018) $(v,C)$ $17.56$ $14.40$ $17.56$ $14.40$
SE(3)-Tr. (Fuchs et al., 2020) $(v,C)$ $21.67$ $16.77$ $21.67$ $16.77$
Lie Neurons (Lin et al., 2024a) $(v,C)$ $16.86$ $13.65$ $16.86$ $13.65$
@@
ReLN (Ours) $v$ $16.85$ $12.70$ $16.85$ $12.70$
ReLN (Ours) $(v,C)$ $16.49$ $13.02$ $16.49$ $13.02$
ReLN (Ours) $(v,\log C)$ 13.92 11.04 13.92 11.04
ReLN (Ours) (no semisimple) $(v,\log C)$ $16.27$ $12.65$ $16.27$ $12.65$